# Augmenting NER Datasets with LLMs: Towards Automated and Refined Annotation

### v1

@@ Abstract
In the field of Natural Language Processing (NLP), Named Entity Recognition (NER) is recognized as a critical technology, employed across a wide array of applications. Traditional methodologies for annotating datasets for NER models are challenged by high costs and variations in dataset quality. This research introduces a novel hybrid annotation approach that synergizes human effort with the capabilities of Large Language Models (LLMs). This approach not only aims to ameliorate the noise inherent in manual annotations, such as omissions, thereby enhancing the performance of NER models, but also achieves this in a cost-effective manner. Additionally, by employing a label mixing strategy, it addresses the issue of class imbalance encountered in LLM-based annotations. Through an analysis across multiple datasets, this method has been consistently shown to provide superior performance compared to traditional annotation methods, even under constrained budget conditions. This study illuminates the potential of leveraging LLMs to improve dataset quality, introduces a novel technique to mitigate class imbalances, and demonstrates the feasibility of achieving high-performance NER in a cost-effective way.
@@ 1 Introduction
In the field of Natural Language Processing (NLP), Named Entity Recognition (NER) is a critical task that involves identifying and classifying named entities in a given text into predefined categories, including persons, organizations, locations, temporal expressions, quantities, monetary values, and percentages (Mohit, 2014). NER is integral to a wide range of applications, from basic information retrieval and content classification to more complex tasks like question answering and machine translation (Guo et al., 2009; Aliod et al., 2006; Babych & Hartley, 2003). The enhancement of NER model performance not only directly contributes to the improvement of broad NLP applications but also plays a crucial role in enhancing our understanding and engagement with text-based data (Cheng & Erk, 2019; Etzioni et al., 2005).
In recent years, NER methods leveraging the in-context learning capabilities of Large Language Models (LLMs) have been proposed, demonstrating high performance close to that of models specialized in NER by carefully selecting the examples provided to the LLMs (Wang et al., 2023). This indicates the effectiveness of utilizing LLMs’ advanced text comprehension abilities for NER tasks. On the other hand, inference with LLMs requires significant computational resources and involves lengthy latency before obtaining inference results, making them unsuitable for applications requiring the rapid processing of huge amounts of text data. Therefore, we believe it is crucial to develop approaches that leverage the language understanding capabilities of LLMs while requiring fewer computational resources and achieving faster NER task processing. Consequently, we focus on improving the quality of datasets used for training NER models with the aid of LLMs, aiming to enhance the performance of smaller, faster, dedicated NER models.
The performance and effectiveness of NER models rely heavily on the quality of annotations in the datasets used during their training (Mitkov et al., 2000). Currently, the majority of datasets employed for training NER models are annotated manually by humans. This manual annotation process is fraught with several challenges, such as inconsistent quality of annotations, the presence of missed and incorrect annotations, significant time and financial costs, and the need for ethical considerations when dealing with text data that may contain offensive content, among others (Katiyar & Cardie, 2018; Fort et al., 2011).
In this study, we propose a novel approach to address these issues, particularly the problems of missed annotations and the time and financial costs associated with manual annotation, intending to enhance the quality of datasets and thereby strengthen the performance and effectiveness of NER models. Our proposed method leverages LLMs to supplement the missed annotations in datasets that have been manually annotated (where the presence of missed annotations is assumed) as shown in Figure 1. By utilizing a hybrid approach that combines manual annotations with annotations generated by large LLMs, we aim to recover from the noise introduced by missed annotations in the dataset, thereby enhancing the quality of the datasets used for training NER models at a low cost and in an automated manner.
@@
In our proposed method, a detailed analysis of the annotations by LLMs revealed that a single expression could be assigned multiple named entity labels. Also, an imbalance in the data volume per label exists, which in turn causes performance disparities among labels in the NER model. This is particularly critical in the “Miscellaneous” category, where such inequalities are most pronounced. To address these issues, we implemented a label mixing technique for the multiple assigned named entity labels, thereby ensuring the robustness of the NER model by mixing these labels.
In the experiments, the CoNLL03 (Tjong Kim Sang & De Meulder, 2003) and WikiGold (Balasuriya et al., 2009) datasets were enhanced using our proposed hybrid annotation method combining manual human efforts and LLMs, and the improved datasets were then used to train a BERT-based (Devlin et al., 2019) NER specialized model. The experimental results demonstrated that our proposed approach could achieve high performance in training NER models with datasets containing noise such as missed annotations, despite requiring minimal financial cost.
The contributions of this study can be summarized in three main points:
  • We propose a hybrid annotation method combining manual human efforts and LLMs for NER datasets, automatically recovering from noise such as missed annotations contained within the datasets (Figure 2).
  • By incurring only a minimal financial cost, we enhance the performance of NER-specialized models (Table 1).
  • In response to the issue of imbalance among named entity labels arising from LLM-based annotations, we propose a label mixing technique, ensuring the robustness of the NER model (Figure 4).
2 Related Works
2.1 Reliability of NER datasets
@@
The reliability of datasets used in NER tasks has been scrutinized. Studies, such as the one presented in Han et al. (2023a), highlight the inherent errors in these datasets. These inaccuracies range from mislabeled entities to inconsistent categorization, impacting the overall performance and reliability of NER systems (Mishra et al., 2020; Mitkov et al., 2000).
2.2 LLMs for NER tasks
Wang et al. (2023) attempted to utilize GPT, a large-scale language model (LLM) (Brown et al., 2020; Ouyang et al., 2022), for annotation tasks and validated its effectiveness. Additionally, they tested the output format and improved annotation accuracy using few-shot learning. LLMs’ advantages for NER tasks are high consistency and efficiency. Conventional methods trained with manual annotation tend to vary due to annotator subjectivity and fatigue. In addition, the annotation cost for LLMs is much cheaper than that of human annotators. The disadvantage of their work is that the computational cost of inference is very high each time. While their approach directly solves the NER task, which differs, our approach includes augmenting and creating the NER dataset from scratch.
2.3 Challenges and Remedies in Automatic Annotation
@@
There are several challenges in automatic annotation. First, Ghosh et al. (2023) proposes ACLM (Attention-map aware keyword selection for Conditional Language Model fine-tuning), an attention map-based data expansion method to address the problem of data imbalance. It should be noted, however, that this method does not take advantage of LLMs. Next, Shen et al. (2023b) proposes using multiple prompt templates to address the problem of multiple label assignments to the same entity. Furthermore, DiffusionNER in Shen et al. (2023a) and Zhang et al. (2022) address the issues of noise and generative bias by utilizing causality-based methods. To address these issues, we propose a hybrid approach that combines manual annotation with automatic LLMs annotation.
Other related studies in NER that must be mentioned are Li et al. (2020) for a comprehensive review of NER deep learning methods, Liu et al. (2021) addresses noise from external knowledge bases, and Wang et al. (2022) propose MINER (Mutual Information based Named Entity Recognition) for out-of-vocabulary support.
While the above diverse approaches address the NER challenge, this study proposes a unique methodology that focuses on improving the reliability of the dataset itself and leads to better accuracy for NER tasks.
3 Methods
LLM-based Annotation
In linguistic data automation using LLM, the GPT-NER methodology, as delineated in Wang et al. (2023), emerges as a paradigmatic example. This method profoundly influences our study, harnessing LLMs to facilitate automated annotation. Central to the GPT-NER strategy is the LLM’s capability to discern and earmark specific entities within the text, encapsulating these segments with distinct character strings, specifically “@@” and “##”. For illustrative purposes, given an input such as “I live in Tokyo”, wherein the objective is to pinpoint LOC (location) entities, the LLM’s output would be “I live in @@Tokyo##”. This delineation effectively signifies the accurate detection of “Tokyo” as a location entity, with the “@@” and “##” serving as the entity’s boundaries.
Additionally, this approach empowers the LLM to apprehend both the intrinsic properties of entities and their interplay with the input-output schema, achieved through the integration of few-shot examples. In alignment with this methodology, we have formulated the ensuing prompt structure. {itembox}[l]Instructions for the Large Language Model You are an excellent linguist.
@@ 3 Methods
Within this structured prompt design, the [target entity type] slot accommodates various entity types, including but not limited to PER, ORG, LOC, and MISC. Concurrently, the [definition target entity type] slot is allocated for the explicit definition of these enumerated entity types. Moreover, the [Example $K$ Input] and [Example $K$ Output] slots are respectively reserved for the input text and its corresponding output text, collectively functioning as few-shot examples.
As this prompt dictates, the entity recognition process is methodically applied to each target entity type within the text targeted for annotation. In scenarios where this strategy involves multiple iterations of annotation for each type of entity, a single word may be assigned multiple different entity types. To resolve instances of such overlapping entities, the simplest method we adopt is to select one entity from those multiple assignments randomly.
The criterion for selecting few-shot examples is inspired by the GPT-NER approach, entailing the selection of examples from validation data. This selection is predicated on the cosine similarity between token embeddings present in the input text.
Hybrid: LLM-Based and Manual Annotation
To enhance the quality of the NER dataset, we created Hybrid annotations by merging annotations from LLMs with those generated manually. This merging process involved carefully analyzing the annotations from both sources, ensuring consistency and accuracy. The merged data combines all entity labels from the manual data and those from LLM annotation that do not overlap the same position as those from the manual data. Since the entity labels of the manual data are expected to be of higher quality than the ones from LLM annotation, they are merged in preference to the ones from LLM annotation. The merged dataset offers a more robust and diverse set of annotations, mitigating biases and errors that might exist in either of the individual sources.
@@
Label Mixing in LLM-Based Annotation
Addressing the issue, as mentioned earlier, of multiple labels being assigned to a single word, we extended the Mixup (Zhang et al., 2018) technique, traditionally used in image processing, to NER tasks. Mixup is a data augmentation method that improves model performance and robustness by mixing two inputs and outputs. Our approach involves blending two different entity labels for a single token to create new annotations. This method is formally represented as follows:
$y_{\text{mix}}=\lambda\times y_{i}+(1-\lambda)\times y_{j}$
Where $y_{i}$ and $y_{j}$ are two labels annotated NER tags, and $\lambda\in[0,1]$ is a parameter determining the mixing ratio. Following the mixup technique (Zhang et al., 2018), $\lambda$ is drawn from the beta distribution ($\alpha=\beta=0.2$). Compared to randomly selecting one label from multiple labels assigned to a single word, this technique contributes to a more nuanced understanding of tokens that could belong to multiple entity types, enhancing the dataset’s richness and classification flexibility.
@@ 4 Experiments
4.1 Setup
In our research, we conducted experiments using two distinct datasets. Firstly, we employed the CoNLL03 dataset (Tjong Kim Sang & De Meulder, 2003), recognized as a standard benchmark in the field of NER. Additionally, acknowledging recent advancements in this field models that have led to a saturation in performance improvements on the CoNLL03 dataset, we also included the WikiGold dataset (Balasuriya et al., 2009). The WikiGold dataset was specifically chosen due to its higher level of challenge and the observation that performance levels achieved on it have not yet paralleled those on the CoNLL03 dataset. This was split into training, validation, and test sets in a 7:1:2 ratio for the experiments.
For the automated annotation process, we leveraged ChatGPT (gpt-3.5-turbo-0613) API [^1] https://chat.openai.com as LLM, and we set the number of few-shot examples at 32.
Finally, to validate the effectiveness of our annotation approach, we fine-tuned a sequence classification model based on a pre-trained BERT (Devlin et al., 2019), utilizing the datasets annotated through this methodology. The evaluation of the NER task is in line with Tjong Kim Sang & De Meulder (2003), and we present weighted average F1 scores in this paper.
A detailed description of the hardware and software configurations is also provided in Appendix appendix A to ensure reproducibility.
4.2 Results
Noise Recovery with LLM-based Annotations:
We assessed the capability of LLM-based annotations to recover from noisy data. Han et al. (2023b) pointed out some annotation errors, such as missing annotations or label switching, in the CoNLL03 dataset (Tjong Kim Sang & De Meulder, 2003). We prepared noised data that reproduces this problem by removing annotations for some entity annotations. We show that it can be improved by combining LLM-based annotation in the previously described manner. We experimented with 20%, 40%, 60%, 80% entity annotations removed. By comparing the noised datasets to those enhanced through our proposed combining methodology, we established a baseline for improvement. Figure 2 presents a comparative analysis of these datasets, highlighting the efficiency of our approach to noise reduction. It was shown that even when noise is severe, it can be significantly improved by combining LLM-based annotations.
@@ 4.2 Results
Comparison Under the Same Budget:
We evaluated our approach under identical budget constraints by drawing parallels with the study of  Ding et al. (2023). Our experiments were conducted on three different budgets. We created composite datasets including different numbers of samples from manual and LLM-based annotations. The middle budget (\$152) is the cost of LLM-based annotation of all training data of the CoNLL03 dataset, and the costs of a quarter and a quadruple of it are assumed for small (\$38) and large (\$608) budgets, respectively. The relationship between the balance of the number of data in each budget and performance is shown in Table 1.
Addressing the Imbalanced Class Issue:
Figure 3 shows that there is a difference in the number of annotations for each entity type. This imbalance issue is even greater with annotation by LLM. Also, the performance difference of LLM-based annotation per entity type is large compared to manual annotation. To tackle the class imbalance issue, we motivated the use of our proposal, label mixing. We present the comparison between LLM-based annotation and label mixing in Figure 4. Label mixing reduced the performance difference between entity types. Therefore, we believe that soft labeling of LLM-based annotations according to possible entities rather than one-hot hard labels will prevent model training in a biased manner.
Figure 4: Improvement of each entity type and micro average with label mixing.
@@
In addition to the analysis above, we further explore the impact of LLM-based annotations and label mixing on the CoNLL03 and WikiGold datasets, particularly focusing on the number of entities that overlap between two classes. Table 2 presents a detailed quantification of these overlapping entities between different entity types when annotated by LLM. The fact that the entities overlapping between classes constitute only 521 cases (2.65%) of the CoNLL03 dataset and 109 cases (6.80%) of the WikiGold dataset, and yet the application of label mixing to these specific data points results in a performance improvement across the dataset, is a suggestive observation. This observation implies that the strategic application of label mixing to even a small subset of data can improve the performance balance between entity types. Moreover, the fact that the number of entities overlapping between classes such as LOC and MISC with ORG is comparatively higher than other combinations, while there is less overlap between PER and MISC, indicates a potential challenge in the discriminatory capability of LLM-based annotations for the former combinations. These observations support the effectiveness of label mixing as a technique to buffer these shortcomings, suggesting that it serves as a valuable strategy for improving the robustness and performance of NER models.
@@ 5 Discussion and Conclusion
Our study extends the use of LLMs beyond annotation tasks. We identified and addressed new NER annotation challenges, proposing novel solutions. Our integration of LLM-based annotations with manual annotation, especially in datasets with noise or limited samples, significantly enhances downstream task performance. This method leverages LLMs’ evolving strengths (Zhao et al., 2023; Min et al., 2023) and introduces label mixing to address biases and imbalances from LLM-based annotations. Our work illuminates LLMs’ role in enhancing NER, underscoring the need for ongoing adaptation and improvement in annotation methodologies.

### v2

# Augmenting and Refining NER Datasets through LLMs

@@ Abstract
Traditional named entity recognition (NER) dataset annotation methods often suffer from high costs and inconsistent quality. This study proposes a hybrid annotation approach that combines human effort with large language models (LLMs) to reduce noise due to the missing labels and improve NER model performance cost-effectively. A label-mixing strategy is also introduced to address class imbalance. Our experiments demonstrate that the proposed method outperforms traditional approaches, even under limited budget constraints.
Augmenting and Refining NER Datasets through LLMs
Yuji Naraki thanks: Equal contributions.    Correspondence to: Yuji Naraki <yuji.1277@gmail.com>, Ryosuke Yamaki <yamaki.ryosuke@em.ci.ritsumei.ac.jp>, Hiroki Naganuma <hiroki.naganuma@proplace.jp>. Independent Researcher                        Ryosuke Yamaki Ritsumeikan University / ProPlace Inc
Yoshikazu Ikeda Osaka University / ProPlace Inc                        Takafumi Horie Ritsumeikan University                        Kotaro Yoshida Science Tokyo
Ryotaro Shimizu ZOZO Research                        Hiroki Naganuma Mila, Université de Montréal / ProPlace Inc
@@ 1 Introduction
Named entity recognition (NER) is a critical task that involves identifying and classifying named entities in a given text into predefined categories (Mohit, 2014). NER models’ performance relies heavily on the quality of annotations in the datasets (Mitkov et al., 2000). Currently, most datasets employed for training NER models are annotated by humans. This manual annotation process is fraught with several challenges, such as annotation errors, time and financial costs (Katiyar and Cardie, 2018; Fort et al., 2011).
@@
This study introduces a novel method to address the challenges of manual annotation by leveraging large language models (LLMs) as annotators. The approach aims to improve dataset quality, which enhances the performance of NER models. Using LLMs for annotation offers several advantages. Human annotators can be subjective and suffer from fatigue, which may reduce accuracy. In contrast, LLMs offer consistency and efficiency throughout the annotation process. Furthermore, while employing LLMs as predictors leads to increasing costs with repeated predictions, employing them as annotators only requires a one-time cost during the annotation process. Afterward, training a predictor based on small models, such as BERT, with the improved dataset can provide adequate performance for the NER task.
Our approach using LLMs as annotators reduces the noise of missed labels by integrating manual and LLM-based annotations (as shown in Figure 1), thus improving the quality of the data set for NER model training in a cost-effective and automatic manner. We find that LLMs tended to assign multiple named entity labels to a single expression. Additionally, an imbalance in data volume across labels leads to performance disparities among labels. To address this, we introduce a label mixing technique that integrates multiple named entity labels, reducing the variance of performance among labels and enhancing the model’s robustness.
Comprehensive evaluation experiments validated the effectiveness of our approach, demonstrating its ability to train NER models on noisy datasets, such as those with missed labels, while requiring minimal financial cost.
2 Related Work
@@
In machine learning, the accuracy and reliability of the dataset used for model training are critical factors directly influencing the performance of the trained model. In this context, research has been conducted to scrutinize the reliability of datasets used for training NER models. Research, such as the one presented in Han et al. (2023a), highlights the inherent errors in these datasets. These inaccuracies range from mislabeled entities to inconsistent categorization, impacting the overall performance and reliability of NER models (Mishra et al., 2020; Mitkov et al., 2000).
@@
There are several challenges in automatic annotation. Ghosh et al. (2023) address the data imbalance problem through an attention map-based data expansion. Shen et al. (2023b) utilize multiple prompt templates to address the problem of multiple label assignments to the same entity. Shen et al. (2023a) and Zhang et al. (2022) address the issues of noise and generative bias by utilizing causality-based methods. To address these problems, we propose a hybrid approach that combines manual and automatic LLM annotations.
3 Methodology
3.1 LLM-based Annotation
In the context of automating linguistic data using LLMs, the GPT-NER method, as described by Wang et al. (2023), provides a key approach. This method plays a significant role in our study by leveraging LLMs for automated annotation. A core aspect of GPT-NER is the ability of the LLM to identify and mark specific entities within text, surrounding these with designated character strings, such as “I live in @@Tokyo##”.
The selection of few-shot examples is based on the approach outlined by Wang et al. (2023), where examples are chosen from validation data using cosine similarity between token embeddings in the input text.
Additionally, this approach enables the LLM to understand both the inherent properties of entities and their relationships within the input-output framework by incorporating few-shot examples. Based on this methodology, we designed the prompt structure as described above.
In this prompt structure, the [target entity type] slot is used for different entity types, such as PER, ORG, LOC, and MISC. The [definition target entity type] slot is allocated for explicitly defining these entity types. The [Example $K$ Input] and [Example $K$ Output] slots are designated for the input text and its corresponding output, serving as few-shot examples.
{itembox}
Instructions for the Large Language Model You are an excellent linguist.
@@ 3.1 LLM-based Annotation
The entity recognition process is systematically applied to each target entity type in the text for annotation. In cases where multiple iterations are required for different entity types, a single word may receive multiple labels. To handle such overlapping labels, we use a simple approach: one label is randomly selected from the multiple assignments.
The few-shot examples are chosen based on the method described by Wang et al. (2023), which involves selecting examples from validation data based on the cosine similarity between token embeddings in the input text.
3.2 Hybrid: LLM-Based and Manual Annotation
To enhance the quality of the NER dataset, we created hybrid annotations by merging labels from LLMs with manually generated ones. This process involved carefully analyzing both sources to ensure consistency and accuracy. Entity labels from the manual data were prioritized, with LLM-based labels added only when they did not overlap. As manual annotations are considered higher quality, they were preferred over LLM-generated ones. The resulting merged dataset offers a more robust and diverse set of labels, reducing biases and errors from either source.
@@ 3.2 Hybrid: LLM-Based and Manual Annotation
3.3 Label Mixing in LLM-Based Annotation
Addressing the issue, as mentioned earlier, of multiple labels being assigned to a single word, we extended the Mixup (Zhang et al., 2018) technique, traditionally used in image processing, to NER tasks. Mixup is a data augmentation method that improves model performance and robustness by mixing two inputs and outputs. Our approach involves blending two different labels for a single token to create new labels. This method is formally represented as follows:
$\boldsymbol{y}_{\text{mix}}=\lambda\times\boldsymbol{y}_{i}+(1-\lambda)\times\boldsymbol{y}_{j}$
Where $\boldsymbol{y}_{i}$ and $\boldsymbol{y}_{j}$ are one-hot encodings of two different labels annotated entity types, and $\lambda\in[0,1]$ is a parameter determining the mixing ratio. Following the mixup technique (Zhang et al., 2018), $\lambda$ is drawn from the beta distribution ($\alpha=\beta=0.2$). Compared to randomly selecting one label from multiple labels assigned to a single word, this technique contributes to a more nuanced understanding of tokens that could belong to multiple entity types, enhancing the dataset’s richness and classification flexibility.
@@ 4 Experiments
Setup. We conducted experiments using two distinct datasets. Firstly, we employed the CoNLL03 (Tjong Kim Sang and De Meulder, 2003), recognized as a standard dataset for NER. Additionally, acknowledging recent advancements in this field that have led to a saturation in performance improvements on the CoNLL03, we also included the WikiGold (Balasuriya et al., 2009). The WikiGold dataset was specifically chosen due to its higher level of challenge and the observation that performance levels achieved on it have not yet paralleled those on the CoNLL03. This was split into training, validation, and test sets in a 7:1:2 ratio for the experiments.
For the automated annotation process, we leveraged ChatGPT (gpt-3.5-turbo-0613) [^1] https://chat.openai.com, and we set the number of few-shot examples at 32.
Finally, to validate the effectiveness of our annotation approach, we fine-tuned a sequence classification model based on a pre-trained BERT (Devlin et al., 2019), utilizing the datasets annotated through this methodology. The evaluation of the NER task is in line with Tjong Kim Sang and De Meulder (2003), and we present weighted average F1 scores in this paper.
A detailed description of the hardware and software configurations is also provided in Appendix A to ensure reproducibility.
Noise Recovery with LLM-based Annotations. We assessed the capability of LLM-based annotations to recover from noisy data. Han et al. (2023b) pointed out some annotation errors, such as missing annotations or label switching, in the CoNLL03 dataset (Tjong Kim Sang and De Meulder, 2003). We prepared noised data that reproduces this problem by removing annotations for some entity. We show that it can be improved by combining LLM-based annotation in the previously described manner. We experimented with 20%, 40%, 60%, 80% entity annotations removed. By comparing the noised datasets to those enhanced through our proposed methodology, we established a baseline for improvement. Figure 2 presents a comparative analysis of these datasets, highlighting the efficiency of our approach to noise reduction. It was shown that even when noise is severe, it can be significantly improved by combining LLM-based annotations.
Comparison Under the Same Budget. We evaluated our approach under identical budget constraints by drawing parallels with the study of Ding et al. (2023). Our experiments were conducted on three different budgets. We created composite datasets including different numbers of samples from manual and LLM-based annotations. The middle budget (\$152) is the cost of LLM-based annotation of all training data of the CoNLL03 dataset, and the costs of a quarter and a quadruple of it are assumed for small (\$38) and large (\$608) budgets, respectively. The relationship between the balance of the number of data in each budget and performance is shown in Table 1.
Addressing the Imbalanced Class Issue. Figure 3 shows that there is a difference in the number of annotations for each entity type. This imbalance issue is even greater with annotation by LLM. Also, the performance difference of LLM-based annotation per entity type is large compared to manual annotation. In terms of the performance difference between labels, the variance of F1-scores when using LLM-based annotation is larger than that of manual annotation. Label mixing reduced the performance difference between entity types. Therefore, we believe that soft labels, according to possible entities, will prevent model training in a biased manner, unlike one-hot labels.
@@
CoNLL03 WikiGold
Annotations mean variance mean variance
LLM-based
w/o label mixing
0.83 0.011 0.63 0.043
LLM-based
w/ label mixing
0.84 0.008 0.63 0.032
Manual 0.89 0.005 0.80 0.009
Table 2: Comparison on the average F1-score and variance between before and after label mixing.
@@
Table 3 presents a detailed quantification of these overlapping entities between different entity types when annotated by LLM. Moreover, the fact that the number of entities overlapping between classes such as LOC and MISC with ORG is comparatively higher than other combinations, while there is less overlap between PER and MISC, indicates a potential challenge in the discriminatory capability of LLM-based annotations for the former combinations. These observations support the effectiveness of label mixing as a technique to buffer these shortcomings, suggesting that it serves as a valuable strategy for improving the robustness and performance of NER models.
@@ 5 Discussion and Conclusion
Our study extends the use of LLMs toward the annotation process. We addressed NER annotation challenges of missing labels and proposed novel solutions. Our integration of LLM-based annotation with manual annotation significantly enhances performance of NER task, especially when using noised datasets or in limited budgets. We have shown that our proposed label mixing address the performance degradation due to biases and imbalances. Our approach capitalizes on the capabilities of LLMs, which are expected to continue advancing in the future (Zhao et al., 2023; Min et al., 2023).